One way to heat a gas is to compress it. A gas at 1 ATM at 25˚C is compressed to 1/10 of its original volume, and it reaches 40 ATM pressure. What is its new temperature?

1,692˚C

919˚C

846˚C

1,500˚C

1,192˚C
Answer #1

Solution:

Using the ideal gas equation,

PV = nRT

PV/T = constant So,

P1V1/T1 = P2V2/T2

T2 = P2V2T1/P1V1

Given, P1 = 1 atm , T1 = 25+273.15 = 298.15 K, P2 = 40 atm , V2 = V1/10

Substitute the values to calculate T2

T2 = 40*[V1/10]*298.15/1*V1

= 1192.6 K

= 919.45oC

Option: B

I hope you understood the problem and got your answers, If yes rate me!! or else comment for a better solutions.